OpenStudy (anonymous):

Evaluate the function f (x)  4 • 7x for x  1 and x = 2. Show your work. i got 4×7^(-1) f(-1) = -28 4×7^2 f(2)=196 but its wrong

Nnesha (nnesha):

type y our question again what's that "box" for ?

OpenStudy (anonymous):

. Evaluate the function f (x) = 4 • 7x for x = -1 and x = 2. Show your work.

Nnesha (nnesha):

okay first replace x by -1 whatdo you get ?

OpenStudy (anonymous):

4×7^(-1) i get this

Nnesha (nnesha):

is it \[\huge\rm f(x) = 4 \times 7^x\] ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

ohh okay remember exponent rule \[\huge\rm x^{-m} = \frac{ 1 }{ x^m }\]

OpenStudy (anonymous):

okay, so do i still replace the x with -1 or do i somehow have to make it a positive 1

OpenStudy (jdoe0001):

she meant \(\bf a^{-{\color{red} n}} \implies \cfrac{1}{a^{\color{red} n}}\qquad \qquad \cfrac{1}{a^{\color{red} n}}\implies a^{-{\color{red} n}} \\ \quad \\ % negative exponential denominator a^{{\color{red} n}} \implies \cfrac{1}{a^{-\color{red} n}} \qquad \qquad \cfrac{1}{a^{-\color{red} n}}\implies \cfrac{1}{\frac{1}{a^{\color{red} n}}}\implies a^{{\color{red} n}} \) so..... yes

OpenStudy (anonymous):

would the answer be 21 or -28

Nnesha (nnesha):

x is base and m is exponent when there is negative exponent you should change that it's reciprocal

Nnesha (nnesha):

to*

Nnesha (nnesha):

21 how 7^(-1) = ???

OpenStudy (jdoe0001):

\(\bf \large { f(x)=4\cdot 7^x\qquad \begin{cases} x=-1\\ x=2 \end{cases} \\ \quad \\ 4\cdot 7^{{\color{red}{ -1}}}\implies 4\cdot \cfrac{1}{7^{\color{red}{ 1}}}\qquad and\qquad 4\cdot 7^{{\color{red}{ 2}}} }\)